Write y=-4x^2+8x-1 in vertex form.

Answer by stanbon(75887)

Write y=-4x^2+8x-1 in vertex form.
y+1-4(1^2)=-4(x^2-2x+1^2)
y-3=-4(x-1)^2
Vertex at:
x=1, y=3
Cheers,
Stan H.